An isosceles triangle has two sides of equal length.The length of the third side is 3 ft less than twice the length of an equal side.Find the length of the third side when the perimeter is 41 ft.
Answer by sravanik2005(14)

Let a,b,c be the lengths of sides
and a,b be equal sides given c=2a-3
perimeter=a+b+c=41
a+a+2a-3=41
4a=44 =>a=11
so,length of third side=2(11)-3=19ft